# Some geometric series for Euler’s constant

### 07 May 2026 v2

@@ Abstract.
Euler’s constant, Euler-Mascheroni constant, asymptotic of sequences
@@ 1. Main result
This gives the sequence $2$, $\frac{7}{3}$, $\frac{8}{3}$, $\frac{133}{45}$, $\frac{16}{5}$, … , $e_{10}=\frac{163287}{40579}$, …. See Table 1 for more values.
Table 1. The coefficients $e_{m}$, $1\leq m\leq 20$
@@ 99 $\frac{20848}{5355}$
As usual, $\gamma$ is Euler constant. Here is the main Theorem.
@@
Let $\ell\geq 2$. Let $(e_{m})$ be the sequence defined by (1). There holds
@@
The level $\ell$ is the number of binary digits used by the integers whose inverse powers are added in Equation (2). In Table 2, we illustrate the numerical values obtained using up to $e_{10}$ for $\ell$ being equal to $2$, 33, …, $7$ (so that the computations could have been done by a diligent human calculator), and also up to $e_{20}$ for $\ell=2,3,4$. The computed approximations to Euler’s constant $\gamma$ in this table are truncated not rounded.
Table 2. Table of approximations of $\gamma$
@@
A numerical implementation using Python and the mpmath library is provided at https://burnolmath.gitlab.io/dyadic-gamma/ (it uses by default $\ell=8$). One can express (see Equation (13)) $e_{m}$ as an explicit finite linear combination of the Bernoulli numbers $B_{k}$, $1\leq k\leq m$, but this uses terms which are much larger than $e_{m}$ itself. So, it seems that using the recurrence (1), which is stable numerically, is more appropriate for computing $e_{m}$’s with hundreds of decimal places. However, an algorithm using (1) appears to have an intrinsic (at least) quadratic cost, and is not suited for large scale computations.
[…]
@@
Figure 1. $e_{m}-(\log 2)^{-1}H_{m+1}$ vs $\log_{2}(m)$, $16\leq m\leq 500$
@@
There holds
$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(\frac{\log m}{m})$ (3)
with
$\phi(t)=-(\log 2)^{-1}\sum_{n\in\mathbb{Z},n\neq 0}\Gamma\bigl(1-\frac{2\pi in}{\log 2}\bigr)\zeta\bigl(1-\frac{2\pi in}{\log 2}\bigr)\mathsf{e}^{2\pi int}.$
[…]
Figure 2. $e_{m}-(\log 2)^{-1}H_{m+1}$ as function of $\log_{2}(m)$, and the graph of $\frac{\gamma}{\log 2}+\frac{1}{2}+\phi(t)$, $16\leq m\leq 1000$
As a corollary to the Theorem, $e_{2m}\approx 1+e_{m}$, a fact which does not seem immediately obvious from Equation 1.
@@
Numerical investigations suggest that the logarithm in the numerator of the error term from Equation (3) can be omitted. However we should remain cautious because we can compute in reasonable time only of the order of ten thousand coefficients. To be more precise, the error term appears to have the shape $\psi(\log_{2}(m))/m+o(1/m)$ with some $1$-periodic function $\psi$ averaging to $\frac{3}{2\log(2)}$.
@@
Remark 3.
We will establish concrete explicit upper and lower bounds for $e_{m}$. They are useful, especially the upper bound, to calibrate numerical implementations.
@@ 2. Geometric series for $\gamma$
In [5], there is an integer parameter $b>1$, which here has been set to $b=2$. There are generalizations of Theorem 1 to general $b>2$, where the analog of the recurrence from Equation (1) has additional factors given by the power sums of the base-$b$ digits.
@@
3. Estimates for the coefficients $e_{m}$
In this section, we establish the next proposition as a corollary to the recurrence from Equation (1).
@@ 3. Estimates for the coefficients $e_{m}$
For $m\geq 2$, there holds
$\frac{H_{m+1}}{\log 2}-\numprint{0.35}<e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.31}\;.$
We first establish a less tight bound. We will then re-use the tools from the proof of the Lemma 1 to prove Proposition 1.
Lemma 1.
@@
$\frac{H_{m+1}-1}{\log 2}\leq e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.161}\,.$
Proof.
Let us define $\delta_{n}=e_{n}-(\log 2)^{-1}H_{n+1}$ for $n\geq 0$. We thus have, for $m\geq 1$:
$(2^{m+1}-2)e_{m}=\left\{\begin{aligned} 2^{m+1}&+(\log 2)^{-1}\sum_{j=1}^{m}\binom{m+1}{j}H_{m-j+1}\\ &+\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}.\end{aligned}\right$. (8)
Now (as is well-known):
$H_{m+1}+\sum_{j=1}^{m}\binom{m+1}{j}H_{m-j+1}$ $=\sum_{k=1}^{m+1}\binom{m+1}{k}H_{k}$
$=\int_{0}^{1}\sum_{k=0}^{m+1}\binom{m+1}{k}\frac{1-t^{k}}{1-t}\mathrm{d}t$
$=\int_{0}^{1}\frac{2^{m+1}-(1+t)^{m+1}}{1-t}\mathrm{d}t$
$=\int_{1}^{2}\frac{2^{m+1}-u^{m+1}}{2-u}\mathrm{d}u$
$=2^{m+1}\int_{\frac{1}{2}}^{1}\frac{1-x^{m+1}}{1-x}\mathrm{d}x$
$=2^{m+1}H_{m+1}-2^{m+1}\int_{0}^{\frac{1}{2}}\frac{1-x^{m+1}}{1-x}\mathrm{d}x$
$=2^{m+1}H_{m+1}-2^{m+1}\log 2+2^{m+1}\underbrace{\int_{0}^{\frac{1}{2}}\frac{x^{m+1}}{1-x}\mathrm{d}x}_{\epsilon_{m}}.$
Hence, from (8),
$(2^{m+1}-2)e_{m}=(2^{m+1}-2)\frac{H_{m+1}}{\log 2}+\frac{H_{m+1}}{\log 2}+2^{m+1}\frac{\epsilon_{m}}{\log 2}+\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}.$
$e_{m}=\frac{H_{m+1}}{\log 2}+\underbrace{\frac{H_{m+1}+2^{m+1}\epsilon_{m}}{(2^{m+1}-2)\log 2}+\frac{\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}}{2^{m+1}-2}}_{\delta_{m}}$ (9)
Suppose now by induction that $\delta_{j}\geq-(\log 2)^{-1}$, for $0\leq j<m$, which is true for $m=1$, as $\delta_{0}=-(\log 2)^{-1}$. The above implies $\delta_{m}\geq-(\log 2)^{-1}$ too, so it is true for all $m$.
To obtain an upper bound, we define
$\mu_{m}=\max_{0\leq j\leq m}\delta_{j}.$
In particular $\mu_{0}=\delta_{0}=-(\log 2)^{-1}\approx\numprint{-1.443}$, and $\mu_{1}=\delta_{1}=2-(\log 2)^{-1}\frac{3}{2}\approx\numprint{-0.164}$. The above implies, for $m\geq 1$:
$\mu_{m}\leq\frac{H_{m+1}+2^{m+1}\epsilon_{m}}{(2^{m+1}-2)\log 2}+\mu_{m-1},$
hence by induction, we obtain, for any given $m$:
$p>m$ $\implies\mu_{p}\leq\mu_{m}+(\log 2)^{-1}\sum_{n=m+1}^{p}\frac{H_{n+1}+2^{n+1}\epsilon_{n}}{2^{n+1}-2},$
thus, using
$\frac{1}{2^{n+1}-2}$ $=\frac{1}{2^{n+1}}+\frac{2}{2^{n+1}(2^{n+1}-2)}\leq\frac{1}{2^{n+1}}+\frac{1}{4^{n}},$
we obtain, for any $p>m\geq 0$:
$\delta_{p}\leq\mu_{m}+(\log 2)^{-1}\sum_{n=m+1}^{\infty}\left(2^{-n-1}H_{n+1}+4^{-n}H_{n+1}+\epsilon_{n}+2^{-n+1}\epsilon_{n}\right).$
Now,
$\sum_{n=m+1}^{\infty}2^{-n-1}H_{n+1}=\frac{-\log(1-\frac{1}{2})}{1-\frac{1}{2}}-\sum_{n=0}^{m}2^{-n-1}H_{n+1},$
and
$\sum_{n=m+1}^{\infty}4^{-n}H_{n+1}=4\frac{-\log(1-\frac{1}{4})}{1-\frac{1}{4}}-\sum_{n=0}^{m}4^{-n}H_{n+1}.$
Evaluating numerically for $m=10$ we obtain respectively:
$(\log 2)^{-1}\sum_{n=11}^{\infty}2^{-n-1}H_{n+1}$ $\approx\numprint{0.002237}$
[…]
We bound trivially $\epsilon_{n}=\int_{0}^{\frac{1}{2}}x^{n+1}(1-x)^{-1}\mathrm{d}x<2^{-n-1}$, hence:
$(\log 2)^{-1}\sum_{n=11}^{\infty}(1+2^{-n+1})\epsilon_{n}$ $<(\log 2)^{-1}(1+2^{-10})\sum_{n=11}^{\infty}2^{-n-1}$
$=(\log 2)^{-1}(2^{-11}+2^{-21})\approx\numprint{0.000705}$
So for $p>10$, $\delta_{p}<\mu_{10}+\numprint{0.003}$. A quick look at Table 3 shows that $\mu_{10}=\delta_{1}=2-(\log 2)^{-1}\frac{3}{2}\approx\numprint{-0.164043}$. Hence $\mu_{10}+\numprint{0.003}<\numprint{-0.161}$. This concludes the proof of Lemma 1. ∎
Proof of Proposition 1.
We continue with the same technique as in the previous proof. To improve the upper bound, we separate out $\delta_{1}$:
$(2^{m+1}-2)^{-1}\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}$ $\leq\frac{m(m+1)}{2(2^{m+1}-2)}\delta_{1}$
$\mkern-100.0mu+\frac{2^{m+1}-2-m(m+1)/2}{2^{m+1}-2}\max\{\delta_{j},0\leq j<m,j\neq 1\}.$
As we know that $\delta_{j}\geq\delta_{0}=-(\log 2)^{-1}$, we can further bound this as
$\leq\frac{m(m+1)}{2(2^{m+1}-2)}\bigl(\underbrace{\delta_{1}+|\delta_{0}|}_{2-\frac{1}{2}(\log 2)^{-1}\approx\numprint{1.279}}\bigr)+\max\{\delta_{j},0\leq j<m,j\neq 1\}.$
We again use $(2^{n+1}-2)^{-1}<2^{-n-1}(1+2^{-n+1})$ for $n\geq 1$. So:
$\sum_{n=21}^{\infty}\frac{n(n+1)}{2(2^{n+1}-2)}$ $<(1+2^{-20})\frac{1}{8}\sum_{n=21}^{\infty}\frac{n(n+1)}{2^{n-1}}$
$=(1+2^{-20})\frac{1}{8}\left(\frac{2}{(1-\frac{1}{2})^{3}}-\sum_{k=0}^{20}\frac{k(k+1)}{2^{k-1}}\right)$
$\approx\numprint{0.000121117}$
The product of this with $\delta_{1}-\delta_{0}$ is less than $\numprint{0.0002}$. We still need to take into account the contribution to the upper bound from the series with the $H_{n+1}$’s and the $\epsilon_{n}$’s. We expect something about one thousand times smaller than what we computed in the previous proof for the remainders, and indeed, one obtains:
[…]
$(\log 2)^{-1}\sum_{n=21}^{\infty}(1+2^{-n+1})\epsilon_{n}$ $<(\log 2)^{-1}(1+2^{-20})\sum_{n=21}^{\infty}2^{-n-1}$
[…]
As $\delta_{2}\approx\numprint{-0.311608}$ is, as we see from Table 3, the maximum among the $\delta_{m}$’s, for $m\neq 1$ and $0\leq m\leq 20$, we conclude, with some margin, that
$m\geq 2\implies\delta_{m}<\numprint{-0.3114}\;.$
Similarly, regarding the lower bound, we can separate out $\delta_{0}=-(\log 2)^{-1}$, writing
$(2^{p+1}-2)^{-1}\sum_{j=1}^{p}\binom{p+1}{j}\delta_{p-j}\geq\frac{p+1}{(2^{p+1}-2)}\delta_{0}+\min\{\delta_{j},1\leq j<p\},$
where we used the fact that we know $\delta_{j}\leq 0$ for all $j$’s. Now working again with $p>10$, we obtain (as the other contributions involving the $H_{n+1}$’s and the $\epsilon_{n}$’s are positive):
$p>10\implies\delta_{p}\geq\delta_{0}\sum_{n=11}^{\infty}\frac{n+1}{(2^{n+1}-2)}+\underbrace{\min\{\delta_{j},1\leq j\leq 10\}}_{{}=\delta_{3}}.$
Now
$\sum_{n=11}^{\infty}\frac{n+1}{(2^{n+1}-2)}$ $<(1+2^{-10})\sum_{n=11}^{\infty}\frac{n+1}{2^{n+1}}$
$=(1+2^{-10})\frac{1}{2}\left(\frac{1}{(1-\frac{1}{2})^{2}}-\sum_{k=0}^{10}(k+1)2^{-k}\right)$
$\approx\numprint{0.006354}.$
The product of this with $\delta_{0}$ is about $\numprint{-0.009167}>\numprint{-0.01}$ and we thus can conclude that for all $m\geq 1$, $\delta_{m}>\numprint{-0.01}+\delta_{3}>\numprint{-0.35}$. ∎
4. The asymptotic behavior of $(e_{m})$
@@
$\Re s>1\implies u_{m}(s)=\sum_{l=0}^{\infty}\biggl(\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-ls}\,.$ (10)
@@
$0\leq\int_{0}^{1}t^{m}\mathrm{d}t-2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\leq\frac{1}{2}2^{-l}.$ (11)
So, at first for $\Re s=1$:
$u_{m}(s)$ $=(m+1)^{-1}\sum_{l=0}^{\infty}2^{-l(s-1)}-\sum_{l=0}^{\infty}\biggl((m+1)^{-1}-2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
$(m+1)u_{m}(s)$ $=\frac{2^{s}}{2^{s}-2}-\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
$(m+1)c_{m}(s)$ $=1-\frac{2^{s}-2}{2^{s}}\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}.$
Thanks to estimate (11), this gives the analytic continution to $\Re s>0$, and we can now compute the value of the derivative at $s=1$:
$(m+1)c_{m}^{\prime}(1)=-\log(2)\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr).$
@@
$e_{m}=\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)\,.$ (12)

### 15 May 2026 v4

@@ Abstract.
Euler’s constant, asymptotics, Riemann zeta function, Bernoulli numbers and polynomials, linear recurrences involving binomial coefficients
@@ 1. Main result
This gives the sequence 0, 22, $\frac{7}{3}$, $\frac{8}{3}$, $\frac{133}{45}$, $\frac{16}{5}$, … , $e_{10}=\frac{163287}{40579}$, …. See Table 1 for more values.
Table 1. The coefficients $e_{m}$, $1\leq m\leq 20$.
See also https://oeis.org/A372422.
@@ 99 $\frac{20848}{5355}$
They can be expressed, as we will see, in terms of Bernoulli numbers (using the convention $B_{1}=-\frac{1}{2}$):
$e_{m}=-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{2^{k}}{2^{k}-1}=1-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{1}{2^{k}-1}\,.$ (2)
(the second formulation supposes $m\geq 1$).
Let $\gamma$ be as usual the Euler-Mascheroni constant. Here is the main Theorem:
@@
Let $\ell\geq 2$. Let $(e_{m})$ be the sequence defined by Equation (1). There holds
@@
The level $\ell$ is the number of binary digits used by the integers whose inverse powers are added in Equation (3). In Table 2, we illustrate the result of computations well within the reach of a diligent human calculator: they use each $10$ terms of the series for $\ell$ respectively equal to $2$, 33, …, $7$, and $20$ terms for $\ell=2,3,4$.
Table 2. Table of approximations of $\gamma$ (decimal expansions in the middle column are truncated)
@@
A numerical implementation using Python and the mpmath library is provided at https://burnolmath.gitlab.io/dyadic-gamma/ (it uses by default $\ell=8$). The coefficients $e_{m}$ are computed using the recurrence (1) (the larger the $m$, the less the required precision), which induces a quadratic cost which is redhibitory for large scale computations. Equation (2) is not immediately usable numerically for large $m$, as individual terms are much larger than $e_{m}$. Indeed it turns out that $e_{m}\sim\log_{2}(m)$, more precisely the difference is bounded. See Table 3 and Figure 1, which displays intriguing oscillations of $e_{m}-H_{m+1}/\log(2)$ ($H_{n}=\sum_{1\leq j\leq n}j^{-1}$) as a function of $\log_{2}(m)$.
@@
Figure 1. $e_{m}-(\log 2)^{-1}H_{m+1}$ as function of $\log_{2}(m)$, $16\leq m\leq 500$. Notice the oscillations as function of $\log_{2}(m)$ and their small amplitudes.
This is the time to reveal that the sequence $(e_{m})$ (and its peculiar asymptotic) has long been known in the field of theoretical computer science! The numerators and denominators have dedicated OEIS pages A372422 and A372423. Paraphrasing their description in a less precise way, $e_{m}$ is the expected depth of trees related to the process of recursively randomly eliminating people via coin tosses, from a group of initial cardinality $m+1$. It is one among a family of related sequences arising in the theoretical analysis of search algorithms, to which Knuth dedicated an entire chapter of his influential treatise ([14, §6.3 “Digital searching”]). See also the article [9] by Flajolet and Sedgewick which contains a discussion of certain techniques for asymptotic analysis involving contour integrals. We were led to the sequence $(e_{m})$ via the analysis of zeta series with missing digits (inclusive of the case with no missing digits…) [4, 5], so the occurrence in the context of “digital” search algorithm is in retrospect perhaps not so surprising.
The quantity $e_{m}$ is $d_{m+1}$ from [16, §4]. In this reference, Prodinger establishes an asymptotic $d_{m}\sim\log_{2}(m)+\frac{1}{2}-\delta_{2}(\log_{2}(m))$ ([16, Thm. 14]) where $\delta_{2}$ is a $1$-periodic function of zero average and small amplitude, given as a Fourier series involving the values of $\Gamma(s)\zeta(s)$ on the line $\Re(s)=1$. The order of the error is not indicated in [16]. Earlier, Szpankowski published a general theorem about binomial sums with Bernoulli coefficients of the type of Equation (2). This Theorem 4.1 of [17] claims a general error bound which would translate here in our notation into $O(m^{-1})$. Unfortunately, this proof has a gap. [^1] In the study of $\delta_{1}(n,s)$ of [17, eq. (4.12)], the estimate $\binom{n+1}{k}n^{-k}=k!^{-1}\bigl(1+O(n^{-1})\bigr)$ is used for $1\leq k\leq n$ but is of course valid only for bounded $k$’s, or at least under some restriction on $k$ relative to $n$. Hence, the proof given for the total contribution for $j>\log_{d}(n)$ is invalid. We shall explain, in our context, how to fix it.
Knuth had studied earlier in his treatise [14] similar such combinations of Bernoulli numbers. A particular example is the topic of [14, Ex. 6.3-34] (which is not about the sequence $e_{m}$, but a related one). There, an asymptotic formula is obtained with a claimed $O(n^{-1})$ bound for the error. But the proof given ([14, page 727]) is not valid. [^2] The exercise [14, 6.4-34, p. 510] (page 503 in the first edition) tells the reader to prove that a series $\sum_{j}\bigl(S_{n-1}(2^{j})-T_{n}(2^{j})\bigr)$ is $O(n^{-1})$ (using here notation from the answer). The answer on page 727 however appears to be about $\sum_{j}\bigl(S_{n}(2^{j})-T_{n}(2^{j})\bigr)$. But this latter series badly diverges, its general term goes to -$\infty$! Besides, the claim $S_{n}(m)=T_{n}(m)+O(\mathsf{e}^{-n/m}n/m^{2})$ is wrong (it is correct if $m/n$ is bounded). The $O(n^{-1})$ bound itself though is correct, as we shall establish. We shall give a proof for this $O(n^{-1})$.
We give a detailed self-contained approach for the $(e_{m})$ asymptotic.
@@
The sequence $(e_{m})$ obeys the following asymptotic:
$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(m^{-1}),$ (4)
where $\phi$ is a $1$-periodic function, which is analytic for $|\Im t|<\pi/(2\log(2)$), has zero mean, and verifies
$\phi(t)=-t-\frac{1}{2}+\sum_{l=0}^{\infty}\bigl(1-\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\bigr)-\sum_{l=-1}^{-\infty}\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\,$, (5)
and
$\phi(t)=-(\log 2)^{-1}\sum_{n\in\mathbb{Z},n\neq 0}\Gamma\bigl(1-\frac{2\pi in}{\log 2}\bigr)\zeta\bigl(1-\frac{2\pi in}{\log 2}\bigr)\mathsf{e}^{2\pi int}\,.$ (6)
Figure 2. $e_{m}-(\log 2)^{-1}H_{m+1}$ as function of $\log_{2}(m)$, and the graph of $\frac{\gamma}{\log 2}+\frac{1}{2}+\phi(t)$, $16\leq m\leq 1000$, $t=\log_{2}(m)$.
@@
Due to the $O(m^{-1})$ error term it does not matter here whether we are talking about $e_{m}$ ($d_{m+1}$) or about $d_{m}$ from [16, Thm. 14].
Remark 2.
Numerical computations suggest that the $O(m^{-1})$ error term in (4) is $m^{-1}\psi(\log_{2}(m))+o(m^{-1})$ with some $1$-periodic function $\psi$ averaging to $\frac{3}{2\log(2)}$ (whether we use $m$ or $m+1$ now matters for what $\psi$ actually is). Related to this, for $m$ up to a few thousands, $\frac{H_{m+1}}{\log(2)}-\frac{\gamma}{\log 2}+\frac{1}{2}$, which differs from $\log_{2}(m)+\frac{1}{2}$ by a term equivalent to $\frac{3}{2\log(2)m}$, proves to be a much better approximation to $e_{m}$ than $\log_{2}(m)+\frac{1}{2}$.
Remark 3.
We shall obtain the Fourier series (6) as a corollary to (5), which itself will be derived by elementary means. This is thus quite different from the contour integral approach sketched by Prodinger in [16] and Szpankowski [17]. On discovering the gaps regarding error bounds, we expected that the methods of complex analysis, for which an account is to be found in [9] by Flajolet and Sedgewick (see also [11] and the book [10] for many resources), would be the tools of choice. As it turns out, we obtain here Equation (4) with its $O(m^{-1})$ error bound with no recourse to them. It is probable though that they could prove of use for refinements.
@@ 2. Geometric series for $\gamma$
In [5], there is more generally an integer parameter $b>1$, (which would be $M$ in the notation of [14, §6.3]) which in Equation (7) has been set to $b=2$. And there are generalizations of Theorem 1 to general $b>2$, where the analog of the recurrence from Equation (1) has additional factors given by the power sums of the base-$b$ digits. Already in the Treatise of Knuth [14, §6.3], such equations arise, with an integer $M>1$ like the $b$ of [5].
@@
3. Elementary bounds
In this section, we mention two elementary estimates which can be established using only the recurrence (1). As such results are superseded for large $m$ (if one goes through the proof to get explicit bounds) by Theorem 2 and by explicit numerical results for $m$ of moderate size, we omit the proofs and refer the interested reader to https://arxiv.org/abs/2603.29998v1.
@@ 3. Elementary bounds
$\frac{H_{m+1}-1}{\log 2}\leq e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.161}\,.$ (11)
More precisely, for $m\geq 2$:
$\frac{H_{m+1}}{\log 2}-\numprint{0.35}<e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.31}\;.$ (12)
4. The coefficients $e_{m}$ via power sums
@@
$\Re s>1\implies u_{m}(s)=\sum_{l=0}^{\infty}\biggl(\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)2^{-ls}\,.$ (13)
@@
$0\leq\int_{0}^{1}t^{m}\mathrm{d}t-2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\leq\frac{1}{2}2^{-l}.$ (14)
So, at first for $\Re s>1$:
$u_{m}(s)$ $=(m+1)^{-1}\sum_{l=0}^{\infty}2^{-l(s-1)}-\sum_{l=0}^{\infty}\biggl((m+1)^{-1}-2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
$(m+1)u_{m}(s)$ $=\frac{2^{s}}{2^{s}-2}-\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
$(m+1)c_{m}(s)$ $=1-\frac{2^{s}-2}{2^{s}}\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}.$
Thanks to estimate (14), this gives the analytic continuation to $\Re s>0$, and we can now compute the value of the derivative at $s=1$:
$(m+1)c_{m}^{\prime}(1)=-\log(2)\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr).$
@@
$e_{m}=\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq k<2^{l}}\bigl(\frac{k}{2^{l}}\bigr)^{m}\biggr)\,.$ (15)
@@
To obtain Equation (16), we use the properties of Bernoulli numbers and polynomials as summarized below, and $2^{k}/(2^{k}-1)$ is obtained as $\sum_{l=0}^{\infty}2^{-kl}$. As $B_{m+1}(1)-B_{m+1}(0)=0$ if $m>0$, the $l=0$ term contributes +1+1 to $e_{m}$, and there also holds:
$m\geq 1\implies e_{m}=1-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{1}{2^{k}-1}\,.$
The Bernoulli polynomials and numbers used in (16) are defined by
$\frac{t\mathsf{e}^{xt}}{\mathsf{e}^{t}-1}=\sum_{n=0}^{\infty}B_{n}(x)\frac{t^{n}}{n!}.$
They verify $B_{n}(x+1)-B_{n}(x)=nx^{n-1}$, hence
$\sum_{0\leq k<n}k^{m}=\frac{B_{m+1}(n)-B_{m+1}(0)}{m+1}\,.$
Also
$B_{m+1}(x)=\sum_{k=0}^{m+1}\binom{m+1}{k}B_{k}x^{m+1-k}\,$,
and $B_{0}(x)=1$, $B_{1}(x)=x-\frac{1}{2}$, $B_{n}=B_{n}(0)$, $B_{2p}=(-1)^{p-1}|B_{2p}|$ ($p\geq 1$), $B_{2p+1}=0$ ($p\geq 1$). That $B_{n}$ is used both for the polynomial and its value at zero should not cause confusion.